Coastal’s current enrollment is 5144 students. This represents a 12% increase from enrollment 5 years ago. What was the enrollment 5 years ago? Round to nearest whole number
Answer by likaaka(51)

Set your equation as so,
5144/x=1+12%
5144/x=112%, in other words 5144 is 112% of what
Convert percentage to a decimal fraction
5144/x=1.12
5144=1.12x, solve for x
5144/1.12=x
x=4592.857142all repeated
Enrollment 5 yrs ago was 4593 students